# Uncovering Grounding IDs: How External Cues Shape Multi-Modal Binding

### v2

@@ A.6.2 POPE: Probing Object-Presence Hallucination
To ensure comparability with CHAIR, we evaluate POPE using questions corresponding to the same 200 randomly selected images from the MSCOCO validation set. For each model response, we compute standard classification metrics: precision, recall, F1 score, and accuracy, defined as follows:
@@ A.6.3 Evaluation on POPE Benchmark
In addition to CHAIR, we evaluate our structured scaffolding method using the POPE benchmark, which probes object hallucination through structured binary questions. As described in Appendix A.6, POPE defines three subsets—Random, Adversarial, and Popular—each targeting different hallucination biases such as language priors, scene co-occurrence, and object frequency.
@@
We use the same 200 MSCOCO validation images as in the CHAIR evaluation. From the POPE benchmark, we select yes/no question subsets corresponding to these images. Each image has six questions per subset (Random, Adversarial, Popular), yielding 1,200 questions per subset and 3,600 in total per model. We evaluate both baseline and structured versions of LLaVA-1.5 and Qwen2.5-VL using standard binary classification metrics: precision, recall, F1 score, and accuracy.
@@
LLaVA-1.5 Random Base 87.25 97.45 76.50 85.71 39.25
OPERA 86.17 96.97 74.67 84.37 38.50
VCD 85.42 93.81 75.83 83.87 40.42
SPARC 85.08 98.85 71.16 82.74 50.25
Ours 89.25 95.20 82.67 88.49 43.42
Popular Base 85.92 94.25 76.50 84.45 40.58
OPERA 84.92 93.92 74.67 83.19 39.75
VCD 83.67 89.82 75.83 82.28 42.17
SPARC 84.67 97.06 71.50 82.34 36.83
Ours 86.08 88.73 82.67 85.59 46.58
Adversarial Base 81.42 86.04 75.00 80.14 43.58
OPERA 82.08 87.67 74.67 80.65 42.58
VCD 80.58 83.79 75.83 79.62 45.25
SPARC 84.92 98.84 70.72 82.46 50.08
Ours 83.50 88.95 76.50 82.26 43.00
@@
Qwen2.5-VL Random Base 89.75 98.38 80.83 88.75 41.08
OPERA 88.83 98.74 78.67 87.57 39.83
VCD 89.91 98.58 81.00 88.93 41.08
SPARC 85.33 95.85 71.50 82.98 36.17
Ours 85.67 98.20 72.67 83.52 37.00
Popular Base 88.83 96.05 81.00 87.88 42.17
OPERA 87.41 96.67 77.50 86.03 40.08
VCD 89.00 96.06 81.33 88.09 42.33
SPARC 84.75 96.85 71.83 82.49 37.83
Ours 85.08 96.88 72.50 82.94 37.42
Adversarial Base 87.42 93.09 80.93 86.53 43.42
OPERA 86.42 94.32 77.50 85.09 41.08
VCD 87.41 92.60 81.33 86.60 43.92
SPARC 81.17 80.25 82.67 81.44 51.50
Ours 83.75 93.55 72.50 81.69 38.75
@@
A.7 Activation swapping (interchange intervention)
@@
A.7.1 Disjoint-symbol swap (no-visual-cue prompt)
To further assess the role of abstract identifiers, we repeat the activation-swap experiment with disjoint symbol sets across contexts. The source image uses {&,\$,#,@}, while the target image uses {+, $\times$, %, !}, with no overlap (Fig. 15). After object-only activation patching from the source into the target (pad =1=1; symbol tokens never swapped), we intentionally query the target using a source symbol (e.g., &), which is not physically present in the target image. The query omits any enumeration of the available symbols so as not to contradict the target’s visual glyphs.
@@ A.7.1 Disjoint-symbol swap (no-visual-cue prompt)
(a) Source: symbols &,\$,#,@.
(b) Target: symbols $\times$,+.
Figure 15: Disjoint-symbol control for the activation-swap test. Object-image activations are transferred with pad =1=1; symbol tokens are never swapped. The target is queried with a source symbol that does not appear in the image, and the prompt allows the answer none to avoid bias.

### v3

@@ A.6.2 POPE: Probing Object-Presence Hallucination
To ensure comparability with CHAIR, we evaluate POPE using questions corresponding to the same 500 randomly selected images from the MSCOCO validation set. For each model response, we compute standard classification metrics: precision, recall, F1 score, and accuracy, defined as follows:
@@ A.6.3 Evaluation on POPE Benchmark
In addition to CHAIR, we evaluate our structured scaffolding method using the POPE benchmark, which probes object hallucination through structured binary questions. As described in Appendix A.6.2, POPE defines three subsets—Random, Adversarial, and Popular—each targeting different hallucination biases such as language priors, scene co-occurrence, and object frequency.
@@
We use the same 500 MSCOCO validation images as in the CHAIR evaluation. From the POPE benchmark, we select yes/no question subsets corresponding to these images. Each image has six questions per subset (Random, Adversarial, Popular), yielding 3,000 questions per subset and 9,000 in total per model. We evaluate both baseline and structured versions of LLaVA-1.5 and Qwen2.5-VL using standard binary classification metrics: precision, recall, F1 score, and accuracy.
@@
LLaVA-1.5 Random Base 87.27 97.47 76.66 85.82 39.53
OPERA 86.17 97.00 74.88 84.52 38.93
VCD 85.43 93.86 75.99 83.99 40.70
SPARC 85.10 98.90 71.34 82.89 36.50
Structured 89.23 95.19 82.75 88.53 43.67
Popular Base 85.93 94.28 76.63 84.54 40.80
OPERA 84.93 93.99 74.78 83.30 39.97
VCD 83.67 89.89 76.06 82.40 42.53
SPARC 83.67 97.06 71.74 82.50 39.67
Structured 86.07 88.77 82.77 85.67 46.90
Adversarial Base 81.43 86.17 75.15 80.28 43.87
OPERA 82.07 87.76 74.88 80.81 43.03
VCD 80.57 83.94 76.01 79.78 45.67
SPARC 84.93 98.90 70.98 82.64 36.27
Structured 83.50 89.03 76.72 82.42 43.43
@@
Qwen2.5-VL Random Base 89.73 98.39 80.91 88.80 41.37
OPERA 88.83 98.76 78.85 87.69 40.27
VCD 89.90 98.55 81.13 88.99 41.43
SPARC 85.33 95.91 71.76 82.10 35.07
Structured 85.67 98.22 72.90 83.69 37.43
Popular Base 88.83 96.07 81.09 87.95 42.40
OPERA 87.40 96.70 77.63 86.12 40.43
VCD 89.00 96.09 81.44 88.16 42.63
SPARC 84.77 96.88 72.02 82.62 37.37
Structured 85.07 96.92 72.72 83.09 37.87
Adversarial Base 87.43 93.16 81.02 86.66 43.83
OPERA 86.43 94.35 77.62 85.17 41.30
VCD 87.40 92.64 81.49 86.71 44.37
SPARC 81.17 80.35 82.75 81.53 51.73
Structured 83.73 93.60 72.65 81.80 39.07
@@
A.7 Hyperparameters for Hallucination-Mitigation Methods
For all hallucination experiments on MSCOCO (CHAIR and POPE), we use a unified decoding configuration unless otherwise specified: greedy decoding with temperature $T=0.0$, top-$p=1.0$, beam size $1$, and a maximum of $512$ new tokens.
Below we list the exact hyperparameters used for each hallucination-mitigation baseline, following the recommendations of the original papers.
VCD (Visual Contrastive Decoding).
We follow the hyperparameters suggested in the VCD paper. The decoding coefficients are fixed to:
$\alpha=1,\qquad\beta=0.1,\qquad\gamma=0.1.$
For the diffusion-based distortion process, we adopt the recommended number of noise steps, $T=500.$ These settings are used for both LLaVA-1.5 and Qwen2.5-VL.
OPERA.
We use the default configuration recommended in the OPERA paper. The method introduces an over-trust logit penalty and a retrospection-allocation mechanism, with parameters:
$N_{\text{can}}=5,\qquad\sigma=50,\qquad\alpha=1,\qquad r=15.$
These unified settings are applied across all evaluated models, including LLaVA-1.5 and Qwen2.5-VL. We use beam size $5$ following the authors’ implementation.
SPARC.
For SPARC, we follow the recommended hyperparameters from the original paper. Global parameters include:
$\alpha=1.1,\qquad\beta=0.1.$
Attention recalibration is applied across all layers. The token-selection threshold $\tau$ and the attention-extraction layer depend on the model:
  • LLaVA-1.5: $\tau=1.5$, selected layer =20=20.
  • Qwen2.5-VL: $\tau=3.0$, selected layer =18=18.
A.8 Activation swapping (interchange intervention)
@@
A.8.1 Disjoint-symbol swap (no-visual-cue prompt)
To further assess the role of abstract identifiers, we repeat the activation-swap experiment with disjoint symbol sets across contexts. The source image uses {&,\$,#,@}, while the target image uses {+, $\times$, %, !}, with no overlap (Fig. 6(b)). After object-only activation patching from the source into the target (pad =1=1; symbol tokens never swapped), we intentionally query the target using a source symbol (e.g., &), which is not physically present in the target image. The query omits any enumeration of the available symbols so as not to contradict the target’s visual glyphs.
@@ A.8.1 Disjoint-symbol swap (no-visual-cue prompt)
A.9 Extended Experiments on Alternative Models
Experiments in Sections 3 and 4 were originally conducted on Qwen2.5-VL (7B). Here, we present additional experiments on models with different sizes and architectures. Figures 15 and 16 show within- and cross-modality attention analyses for InternVL3.5 (8B) and Qwen2.5-VL (3B), which exhibit the same attention progression as Qwen2.5-VL (7B) (Figure 2).
Figure 15: Attention patterns of InternVL3.5 (8B) under baseline (top) and structured (bottom) inputs for the scene description task. (a) Within-modality visual attention. (b) Cross-modality attention.
Figure 16: Attention patterns of Qwen2.5-VL (3B) under baseline (top) and structured (bottom) inputs for the scene description task. (a) Within-modality visual attention. (b) Cross-modality attention.
We also repeated the activation-swap experiment from Section 4 on InternVL3.5 (8B) and Qwen2.5-VL (3B). The models exhibit the same qualitative behavior as Qwen2.5-VL (7B). For InternVL3.5, swap accuracy is 0.75 for shape and 0.82 for color, well above the random chance level of 0.25 for shape and 0.125 for color. For Qwen2.5-VL (3B), swap accuracy is 0.93 for shape and 0.73 for color, indicating a consistent dependence on Grounding IDs. Figure 17 shows the corresponding log-prob diagrams, which follow the same pattern observed in Qwen2.5-VL (7B) (Figure 4(b)).
Figure 17: Average log probabilities of $c$ and $c^{*}$ over valid row–symbol–object combinations for InternVL3.5 (a) and Qwen2.5-VL 3B (b). Rows and columns correspond to the selected query symbols and their associated objects.
Finally, we repeat the hallucination mitigation experiments on MS-COCO using two additional open-source models of different sizes: Qwen2.5-VL (3B) and InternVL3.5 (8B). Results are reported in Table 7.
Model Method $_{s}\downarrow CHAIR$$_{i}\downarrow$
Qwen2.5-VL 3B Baseline 32.40 8.82
Structured 24.40 7.32
InternVL3.5 8B Baseline 32.40 6.09
Structured 26.40 8.08
Table 7: CHAIR results on 500 MS-COCO samples for Qwen2.5-VL (3B) and InternVL3.5 (8B).
A.10 Clarification of Terminology
Several terms used in the main paper relate to mechanisms that have been discussed in prior work on internal representations in language and vision–language models. For clarity, we provide brief descriptions here.
Symbolic.
We use the term symbolic to refer to the fact that external cues (such as characters or markers) function as explicit and discrete symbols that index different regions of the input. They can be consistently referenced by the model both in text and in internal computation.
Identifier.
An identifier denotes an internal code that the model assigns to all tokens associated with the same cue. This identifier allows the model to bind visual and textual elements referring to the same partition.
Latent.
The identifier is latent because it does not appear explicitly in the visible input. While the external cue is visible, the identifier itself arises in the model’s hidden activations and can be detected through probing and representation analysis.
Vector.
We refer to the identifier as a vector because it is expressed as a direction or component in activation space rather than as a discrete token.
Abstract.
The identifier is abstract because it does not encode perceptual details such as color or shape. Instead, it represents the relational property of membership in a specific partition, independent of the visual content of that partition.